OpenStudy (anonymous):

Find 'p' if 3x^2+2x+3p=0 has real & different roots. . . ANSWER = p<1/9

OpenStudy (anonymous):

I have no idea on solving this question....

OpenStudy (anonymous):

A quadratic has real and different roots if the descriminant is greater than zero: \[\Delta > 0\]

OpenStudy (anonymous):

delta?? u mean discriminant?

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

so how D>0 ??

OpenStudy (anonymous):

What do you mean? Set the discriminant to be greater than zero b^2-4ac > 0

OpenStudy (anonymous):

I mean how is discriminant greater than zero can u explain?

OpenStudy (harsimran_hs4):

if d > 0 then there are real and different roots d = 0 real and equal roots d < 0 imaginary roots

OpenStudy (agent0smith):

Real and different...? I'm guessing you mean real and distinct. But yes you can use the discriminant. When the discriminant is less than zero, it will have complex/imaginary roots, as the quadratic formula takes the square root of the discriminant ( http://www.purplemath.com/modules/quadform.htm ) \[3x^2+2x+3p=0 \]\[ax^2+bx+c = 0\] so here, a=3, b = 2 and c = 3p discriminant is \[b^2-4ac > 0 \]

OpenStudy (anonymous):

If D>0 then sqrt(b^2-4ac) is a real number and roots are real and unequal

OpenStudy (anonymous):

am I right?

OpenStudy (harsimran_hs4):

\[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] try putting different values of d and check out

OpenStudy (harsimran_hs4):

yes you are!!

OpenStudy (anonymous):

as D>0 and D=b^2-4ac so, (2)^2-4*3*3p>0 4-36p>0 4>36p 1/9>0

OpenStudy (anonymous):

is this ok?

OpenStudy (anonymous):

You lost a p at the end

OpenStudy (anonymous):

oh sorry I typed it in hurry

OpenStudy (anonymous):

p<1/9